OpenStudy (anonymous):

A conical water tank with the vertex down has a height of 5 meters and a radius of 2 meters. It contains water to a depth of 2 meters at the center. How much work is required to empty the tank over the top?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

7-y-y= -1 7-2y=-1 -2y=-1-7 -2y=-8 y=4